OpenStudy (anonymous):

If the scale of the blueprint is 1 inch = 2 feet, what are the actual dimensions of Trisha's bed?

OpenStudy (anonymous):

https://app125.studyisland.com/pics/189278num5c.png

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@ahamed.

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

Your mother.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Trisha Paytas.

OpenStudy (anonymous):

I am still confused

OpenStudy (anonymous):

Booaaah.

OpenStudy (anonymous):

I need help on the question. Please help me

OpenStudy (anonymous):

How red?

OpenStudy (anonymous):

I need help on the question

OpenStudy (anonymous):

The gun is where?

OpenStudy (anonymous):

Excuse me this is a question that needs to be answered

OpenStudy (anonymous):

Now you're thinking with the portals.

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

Possibly a new discovery. Amen.

OpenStudy (anonymous):

Help me please

OpenStudy (anonymous):

Cornholio.

pooja195 (pooja195):

Is her bed number 1?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

This is amazing.

pooja195 (pooja195):

Umm seems theres no numbers.....im guessing they want you to use a ruler?

OpenStudy (anonymous):

Lord has spoken in dimension.

OpenStudy (anonymous):

Well it logged me out I have a different question

OpenStudy (anonymous):

To become an entity you must unlock your true self. Your concun will efine that to your brain.

pooja195 (pooja195):

@HappySprout whats the question now?

OpenStudy (anonymous):

Above are two different models of a microchip. If the model of the microchip on the left measures 6 cm tall, how tall is the model of the microchip on the right? https://app125.studyisland.com/userfiles/16044q7p2.gif

pooja195 (pooja195):

whats does it mean when they say right or left? And what are the microchips

OpenStudy (anonymous):

The microchip is here.

OpenStudy (anonymous):

https://app125.studyisland.com/userfiles/16044q7p2.gif https://media.studyisland.com/cgi-bin/mimetex.cgi?1%20 \%20\text{cm}%20\%20=%20\%20\frac{1}{8}%20\%20\text{mm}

OpenStudy (anonymous):

Sorry it is not working

OpenStudy (anonymous):

https://app125.studyisland.com/userfiles/16044q7p2.gif https://media.studyisland.com/cgi-bin/mimetex.cgi?1%20 \%20\text{cm}%20\%20=%20\%20\frac{1}{8}%20\%20\text{mm}

OpenStudy (anonymous):

https://app125.studyisland.com/userfiles/16044q7p2.gif \[1 = 1/8mm\]

OpenStudy (anonymous):

The pictures aren't clear. I would assume this has something to do with scales?

OpenStudy (anonymous):

https://app125.studyisland.com/userfiles/16044q7p1.gif \[1 cm= 1/4 mm\]

OpenStudy (anonymous):

Is that better?

OpenStudy (anonymous):

multiply the right side by 6

OpenStudy (anonymous):

36

OpenStudy (anonymous):

well we have 10 mm = 1/4 mm

OpenStudy (anonymous):

remember same units always helps

OpenStudy (anonymous):

10mm is one of the answers

OpenStudy (anonymous):

10 mm is equal to 1 cm

OpenStudy (anonymous):

so if the chip is 6 cm that means its 60 mm

OpenStudy (anonymous):

Oh ok so it would be 6 cm

OpenStudy (anonymous):

so it should be 60mm = 15mm

OpenStudy (anonymous):

the chip is 1/4 the size

OpenStudy (anonymous):

Ok confused

OpenStudy (anonymous):

you said the dimension is 1cm = 1/4 mm right?

OpenStudy (anonymous):

So it's 1/4 of 15?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok so first change that 1cm into mm

OpenStudy (anonymous):

10mm

OpenStudy (anonymous):

now we have 10mm = 1/4 mm

OpenStudy (anonymous):

Your mother.

OpenStudy (anonymous):

they are the same units...we can work with that easier

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so you said the chip is 6 cm correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so that means we can multiply each side by 6

OpenStudy (anonymous):

So that would be 36

OpenStudy (anonymous):

i'm sorry error

OpenStudy (anonymous):

Where at?

OpenStudy (anonymous):

if the height of the chip is 6 cm that means its 60 mm

OpenStudy (anonymous):

It logged me out again sorry let me open a new question and I'll tag you.

OpenStudy (anonymous):

the other chip is 1/4 that size

OpenStudy (anonymous):

It's not the same question

OpenStudy (anonymous):

therefore 60mm = 1.5mm. that should be the correct size

OpenStudy (anonymous):

@mathman1988

OpenStudy (anonymous):

so is this the one 1cm=1/4mm

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

oh ok the answer is 1.5 mm. is that a choice?

OpenStudy (anonymous):

we got 60mm=1.5 mm

OpenStudy (anonymous):

No it isn't here are the answers: 14 cm 16 cm 12 cm 10 cm

OpenStudy (anonymous):

oh they want it in cm

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok so 1cm = 1/4 mm

OpenStudy (anonymous):

therefore 1cm = 0.025 cm

OpenStudy (anonymous):

Sorry my computer messing up. Yes

OpenStudy (anonymous):

and you said that the chip is 6cm long

OpenStudy (anonymous):

see the problem is that the answers aren't matching the problem. are you sure its not 1cm = 1/4 cm

OpenStudy (anonymous):

Yes I am sure.I had to log out and back in sorry

OpenStudy (anonymous):

I can do a different question if you would like

OpenStudy (anonymous):

the answers just don't match the question. you know 1/4 mm is very very small

OpenStudy (anonymous):

right